State the following as percentages and whether they are w/w, w/v or v/v:

200 mL in 400 mL of solution

0.02 mg in 200 mL of solution

10 g in 250 g of solution
Answer #1

●Given

200ml in 400ml of solution

Here the percentage will be in volume by volume as both the mixture are mentioned in ml (v/v)

To calculate in percentage

=200÷400×100

=1÷2×100

=50%

Ans:50%(v/v)

●Given

0.02mg in ml of solution which means the mixture is in weight (mg)=w in volume(ml)=v. So it has to be mentioned as w/v

=0.02÷200×100

=0.0001×100

=0.01%

Ans:0.01%(w/v)

●Given10g in 250g of solution which means the mixture is in weight (g)=w in weight(g)=w.So it has to be mentioned as w/w

=10÷250×100

=0.04×100

=4%

Ans:4%(w/w)